# Behavioral alignment in social networks

### v1

@@ Results
Figure 6: Equilibrium time as a function of clustering coefficient, heterogeneity, and average path length. The average equilibrium time of different networks is shown in red for coordination games and blue for anti-coordination games. Each box in panels a-f represents the average equilibrium time for $50$ networks, showing the median, quartiles, and outliers. Open and solid dots in g-i correspond to the data in a-c and d-f, respectively. Panel adg, beh, and cfi correspond to $\tau$ values of $0.26$, $0.37$, and $0.49$, respectively. The average equilibrium time increases (decreases) in coordination (anti-coordination) games as the clustering coefficient increases, while it decreases (increases) in coordination (anti-coordination) games as the heterogeneity coefficient increases. g-i, The equilibrium time increases in coordination games and decreases in anti-coordination games as the average path length increases. The simulation is repeated $3\times 10^{7}$ times for each network. Parameter values: network size $N=30$ and average degree $\bar{k}=4$.
@@
Once again, the effects of network structure on equilibrium time can be understood via the basic network metric of average path length. Fig. 6g-i depicts the variation in average equilibrium time as the average path length changes. We analyze all the states of the star (see Fig. S7 and Fig. S8 in Supplementary Information) and the chain network (see Fig. S9 and Fig. S11 in Supplementary Information) of size six. We find the average equilibrium time grows (shrinks) as the average path length increases under coordination (anti-coordination) games.
@@
Figure 7: Coordination games are more robust than anti-coordination games. We assess the robustness of the system by randomly adding a new edge to the equilibrium state and measuring the resulting distance (i.e., the number of strategy changes required to reach the final equilibrium state from the initial equilibrium state.) The red (blue) boxes represent the average number of strategy changes in coordination (anti-coordination) games, and each of them is the average over 20 networks. ab, The average number of strategy switches varies with changes in clustering coefficients. cd, The average number of strategy switches varies with changes in degree heterogeneity. We used a network size of $N=30$ and an average degree of $\bar{k}=4$.
However, in some equilibrium states for coordination games in which two strategies coexist, adding a new edge can drive the system to a monomorphic state (see Fig. S13 in Supplementary Information for an example). Thus, adding new edges to networks in coordination games may result in several individuals switching strategies when their neighbors with $A$ strategy increase and the value of $\lfloor\tau k\rfloor$ remains constant. This is why there are particularly sensitive networks when the clustering coefficient equals $0.35$ in Fig. 7a. Meanwhile, in heterogeneous networks, fewer individuals tend to switch their strategy than they do in homogeneous networks. Hubs, which are often present in heterogeneous networks, tend to maintain their current strategy under small perturbations, and the strategy changes among other nodes cannot influence the hubs, yielding greater robustness.
@@ Discussion
However, exceptions to our findings may arise under specific conditions. In the complete bipartite graph, there are typically only two equilibrium states in both coordination and anti-coordination games (see Section 4 of Supplementary Information). This finding lies in contrast to our previous finding that in networks with low clustering coefficients, there are more equilibrium states in coordination games. The number of equilibrium states can fluctuate dramatically when the behavioral switching threshold changes, especially if its product with the degree of the highest frequency in the network is an integer (see Fig. S15 and Fig. S16 in Supplementary Information for examples). An integer product makes it possible for equality to hold in Eq. 3 and Eq. 4, and such individuals are free to adopt either strategy $A$ or $B$ in an equilibrium, resulting in a greater quantity of equilibrium states overall.
Imitation, along with coordination and anti-coordination, are all significant heuristic decision-making patterns. Imitation has received considerable attention due to the potential of imitation systems to complete complex tasks under simple rules [6, 50, 51, 52]. However, the advantages of coordination and anti-coordination games in this regard are more substantial than those of imitation. In imitation systems, decision-makers need not discern only the behavioral information of the individuals they interact with but also the payoffs resulting from these behaviors. On the other hand, in coordinating and anti-coordinating systems, decision-makers can make decisions based only on the observed behaviors of others. The distinction between fixation time in evolutionary game theory [53, 54, 55] and equilibrium time discussed here lies in their respective focuses: fixation time quantifies the time for a single mutant trait to become fixed in the system, whereas our notion of equilibrium time captures the time required to reach an equilibrium state from a randomly initialized state. In other words, fixation time is a specific measure of a particular property, while equilibrium time assesses the system’s overall relaxation time.

### v2

@@ RESULTS
Once again, the effects of network structure on equilibrium time can be understood via the basic network metric of average path length. Figure 6g-i depicts the variation in average equilibrium time as the average path length changes. We analyze all the states of the star (see Fig. S7 and Fig. S8 in Supplementary Information) and the chain network (see Fig. S9 and Fig. S11 in Supplementary Information) of size six. We find the average equilibrium time grows (shrinks) as the average path length increases under coordination (anti-coordination) games.
@@
Figure 7: Coordination games are more robust than anti-coordination games. We assess the robustness of the system by randomly adding a new edge to the equilibrium state and measuring the resulting distance (i.e., the number of strategy changes required to reach the final equilibrium state from the initial equilibrium state.) The red (blue) boxes represent the average number of strategy changes in coordination (anti-coordination) games, and each of them is the average over 20 networks. (ab), The average number of strategy switches varies with changes in clustering coefficients. (cd), The average number of strategy switches varies with changes in degree heterogeneity. We used a network size of $N=30$ and an average degree of $\bar{k}=4$.
@@
However, in some equilibrium states for coordination games in which two strategies coexist, adding a new edge can drive the system to a monomorphic state (see Fig. S13 in Supplementary Information for an example). Thus, adding new edges to networks in coordination games may result in several individuals switching strategies when their neighbors with $A$ strategy increase and the value of $\lfloor\tau k\rfloor$ remains constant. This is why there are particularly sensitive networks when the clustering coefficient equals $0.30$ in Fig. 7a. Meanwhile, in heterogeneous networks, fewer individuals tend to switch their strategy than they do in homogeneous networks. Hubs, which are often present in heterogeneous networks, tend to maintain their current strategy under small perturbations, and the strategy changes among other nodes cannot influence the hubs, yielding greater robustness.
@@
Real-world networks
Figure 8: Average path length determines the equilibrium time in empirical networks. The equilibrium time of 20 real-world networks for both coordination games (a) and anti-coordination games (b) is shown as a function of the average path length. The equilibrium time of each network is normalized by $N$. The equilibrium time increases in coordination games and decreases in anti-coordination games as the average path length of empirical networks increases.
Here we analyze the system dynamics across 20 empirical networks (see Supplementary Table S4 for details) that span a wide range of real-world contexts, including physical and online social relationships (e.g., school friendships and Facebook ties), workplace and organizational structures (e.g., CEO clubs and employee interactions), small-scale communities (e.g., villages and tribal societies), and animal social systems (e.g., bison and sheep). These networks vary substantially in size (from $N=16$ to $N=39$) and average degree (from $\bar{k}=3.55$ to $\bar{k}=22.25$), reflecting the structural diversity of natural and social systems. In each empirical network, we simulate strategy updates from random initial conditions and record the average number of time steps required to reach equilibrium, normalized by the network size $N$. Figure 8 demonstrates how average path length systematically shapes equilibrium time in empirical networks. We find that equilibrium time increases with average path length in coordination games (Fig. 8a), whereas it decreases in anti-coordination games (Fig. 8b). We calculate the Pearson correlation coefficient $r$, which measures the strength and direction of the linear relationship between two variables. Specifically, $r>0$ means that larger values of one variable tend to be associated with larger values of the other, $r<0$ means that larger values of one tend to be associated with smaller values of the other, and $|r|$ closer to 1 indicates a stronger relationship. To check whether the observed correlation could simply be due to random chance rather than a real link between the variables, we use a permutation test to calculate the $p$-value to assess statistical significance. The resulting trends reveal a significant positive correlation between average path length and equilibrium time in coordination games (r=0.87,p<0.001r=0.87,\ p<0.001), and a significant negative correlation in anti-coordination games (r=0.48,p=0.028r=-0.48,\ p=0.028), corroborating the patterns previously identified in synthetic network models such as random-regular and scale-free networks. Therefore, such patterns are not artifacts of any individual network but reflect a robust principle generalizing across diverse empirical systems. These findings highlight the predictive power of average path length and support its theoretical importance in shaping system dynamics in strategic interactions, both in controlled models and in complex, real-world environments.
@@ DISCUSSION
Our analysis of network properties, including the clustering coefficient, heterogeneity coefficient, and the average path length, applies for general complex networks (e.g., Erdős-Rényi, scale-free, small-world, and empirical networks). However, in highly structured special cases, exceptions to our findings may arise. In the complete bipartite graph, for example, there are typically only two equilibrium states in both coordination and anti-coordination games (see Section 4 of Supplementary Information). This finding lies in contrast to our previous finding that in networks with low clustering coefficients, there are more equilibrium states in coordination games. The number of equilibrium states can fluctuate dramatically when the behavioral switching threshold changes, especially if its product with the degree of the highest frequency in the network is an integer (see Fig. S15 and Fig. S16 in Supplementary Information for examples). An integer product makes it possible for equality to hold in Eq. 8 and Eq. 9, and such individuals are free to adopt either strategy $A$ or $B$ in an equilibrium, resulting in a greater quantity of equilibrium states overall.
Unlike reinforcement learning, our model is grounded in introspection-based dynamics, where individuals choose strategies based on neighbors’ behaviors in the immediately preceding round and aim only to improve the current payoff. Compared to link-prediction studies, which aim to infer missing edges from existing network structures and may treat circle-like or tree-like topologies as prediction targets [58, 59], our study adopts a different perspective. Here, the network is used as a fixed experimental condition to examine how structural properties, particularly the average path length, influence the evolutionary dynamics and equilibrium outcomes of coordination and anti-coordination systems. Unlike prior data-driven simulations of multi-dimensional opinion dissemination for specific events [60], our study provides a rigorous theoretical framework for coordination and anti-coordination dynamics on graphs, and finds average path length can serve as a general structural metric, which hold across a variety of networked systems.
Imitation, along with coordination and anti-coordination, are all significant heuristic decision-making patter ns. Imitation has received considerable attention due to the potential of imitation systems to complete complex tasks under simple rules [8, 61, 62, 63]. However, the advantages of coordination and anti-coordination games in this regard are more substantial than those of imitation. In imitation systems, decision-makers need not discern only the behavioral information of the individuals they interact with but also the payoffs resulting from these behaviors. On the other hand, in coordinating and anti-coordinating systems, decision-makers can make decisions based only on the observed behaviors of others. Compared with aspiration dynamics [64], which also relies on local information, our model uses observed behaviors of others while aspiration dynamics uses internal benchmarks. They also differ in structural sensitivity: aspiration dynamics are robust to network variations, whereas our model exhibits a systematic dependence on network topology, governed by the average path length. The distinction between fixation time in evolutionary game theory [65, 66, 67] and equilibrium time discussed here lies in their respective focuses: fixation time quantifies the time for a single mutant trait to become fixed in the system, whereas our notion of equilibrium time captures the time required to reach an equilibrium state from a randomly initialized state. In other words, fixation time is a specific measure of a particular property, while equilibrium time assesses the system’s overall relaxation time.